OpenStudy (anonymous):

how do we derivative 25In(x+1)???

OpenStudy (anonymous):

is it 25/x+1???????

OpenStudy (anonymous):

Well, why do you think it is what you wrote?

OpenStudy (anonymous):

@wio another one hELP me out plzzz:D

hartnn (hartnn):

to be more precise, it'll be \((25\ln (x+1))' = 25 (\ln (x+1))' = 25 (1/(x+1))(x+1)'=25/(x+1)*1=25/(x+1)\) so, yes, its 25/(x+1) but using chain rule.

hartnn (hartnn):

\((25\ln (x+1))' \\ = 25 (\ln (x+1))' \\ = 25 (1/(x+1))(x+1)'\\ =25/(x+1)*1=25/(x+1)\)

OpenStudy (anonymous):

thanks a lot:D

hartnn (hartnn):

welcome ^_^